OpenStudy (anonymous):

4logx-6log(x+2) turn into single log

OpenStudy (anonymous):

solving part by part \[4\log(x) = \log(x^4)\] and \[6\log(x+2) = \log(x+2)^6\] from the law of log, log(A)-log(B) = log(A/B) So the above equation should be \[\log(x^4/(x+2)^6)\]

OpenStudy (anonymous):

how do you solve it?